Find the quadratic equation whose roots are 1, 2 and passes through the point (0,2). Show your work.
Answer by rothauserc(4718)

(x-1)*(x-2) = 0
x = 1 and x = 2 are the zeros and
f(x) = x^2 -3x +2 is the quadratic equation
we are given point (0,2) and need to check that point with the equation
2 = 0^2 -3*0 +2
2 = 2
our quadratic equation checks with the point